A proton moves perpendicular to a uniform magnetic field B with arrow at a speed of 2.10 107 m/s and experiences an acceleration of 1.80 1013 m/s2 in the positive x-direction when its velocity is in the positive z-direction. Determine the magnitude and direction of the field.

magnitude =
Answer #1

acceleration = net force / mass

= qVB / m

where q = charge

V = velocity

B = magnetic field

m = mass of the proton

1.8*10^13 = 1.6*10^-19*2.1*10^7*B / 1.67*10^-27

B = 0.00894 T

= 8.94 mT

magntiude of magnetic field = 8.94 mT

direction is -ve y-direction using right hand rule